OpenStudy (anonymous):

the mean of three numbers is 25. the second number is four less than twice the first. the third number is two more than four times the first. find the smallest number.

OpenStudy (ravege):

25*3=75 so your numbers will be 11,18, and 46 so, 11+18+46=75; 75/3 = 25 so your numbers are 11 ,18 ,and 46